TheKnightKing8273:

The area of artful at octagon is 35cm2. What is the area of the regular octagon with sides eight times as long? PLEASE I REALLY NEED HELP, I NEED AN ANSWER BY 12AM!!

TheKnightKing8273:

Answer choices A:280cm2 B:1600cm2 C:2240cm2 D:350cm2

justus:

The forumla for a regular octagon is A=2(1+ √2)a^2

TheKnightKing8273:

But it gives me a completely different answer.

justus:

this is the formula not the answer

justus:

Because the area varies as the square of the sides, it will result in an increase of 8^2 or 64 times the original size by a factor of 8 for the sides.

justus:

a = 8 cm The area of an octagon = 2 x (1 + √2) x a² Area = 2 x (1 + 1.414213562) x 8² Area = 2 x (2.414213562) x 64 Area = 4.828427124 x 64 Calculate to get the answer.

jhonyy9:

@justus so i think this is a good job - explication - but in calcule when we use sqrt(2) - is enough to use just 1,41 not so much decimals - my opinion

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @justus so i think this is a good job - explication - but in calcule when we use sqrt(2) - is enough to use just 1,41 not so much decimals - my opinion \(\color{#0cbb34}{\text{End of Quote}}\) Oh thanks, I see what you mean. Thank you for your opinion I will keep it simple next time.

jhonyy9:

yw

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @justus The forumla for a regular octagon is A=2(1+ √2)a^2 \(\color{#0cbb34}{\text{End of Quote}}\) formula you wan write there ....